# MARS: Reinforcing Multi-Agent Reasoning of LLMs through Self-Play in Strategic Games

### v1

@@ Appendix F Full results
MARS (Tic.) 94.40 94.69 87.01 60.00 90.00 89.53 45.45
MARS (Kuhn.) 95.80 94.39 86.61 63.33 92.50 89.65 48.48
MARS (Hanabi) 94.40 94.54 86.22 66.67 95.00 88.98 44.95
MARS (Multi.) 95.20 94.54 86.61 66.67 92.50 89.53 50.00
+ fix opponent
MARS (Tic.) 92.40 95.38 83.07 43.33 80.00 87.11 43.94
MARS (Kuhn.) 90.20 95.83 83.46 33.33 72.50 86.69 42.42
MARS (Hanabi) - - - - - - -
+ w/o turn-level advantage estimator
MARS (Tic.) 91.00 95.75 82.28 40.00 72.50 86.34 42.93
MARS (Kuhn.) 91.80 95.45 82.68 36.67 82.50 86.79 41.92
MARS (Hanabi) 91.20 95.91 83.86 40.00 75.00 87.33 41.92
+ w/o agent-specific advantage normalization
MARS (Tic.) 91.20 95.83 83.46 40.00 80.00 87.14 38.58
MARS (Kuhn.) 91.20 95.53 84.65 36.67 75.00 86.63 43.94
MARS (Hanabi) 92.60 96.06 82.68 33.33 80.00 87.11 43.43

### v2

@@ 2. In each of your turns:
MARSHAL
   Tic-Tac-Toe 94.40 94.69 87.01 60.00 90.00 89.53 45.45
   Kuhn Poker 95.80 94.39 86.61 63.33 92.50 89.65 48.48
   Mini Hanabi 94.40 94.54 86.22 66.67 95.00 88.98 44.95
   Generalist 95.20 94.54 86.61 66.67 92.50 89.53 50.00
MARSHAL (w/ Fixed Opponent)
   Tic-Tac-Toe 92.40 95.38 83.07 43.33 80.00 87.11 43.94
   Kuhn Poker 90.20 95.83 83.46 33.33 72.50 86.69 42.42
   Mini Hanabi - - - - - - -
MARSHAL (w/o Turn-Level Advantage Estimator)
   Tic-Tac-Toe 91.00 95.75 82.28 40.00 72.50 86.34 42.93
   Kuhn Poker 91.80 95.45 82.68 36.67 82.50 86.79 41.92
   Mini Hanabi 91.20 95.91 83.86 40.00 75.00 87.33 41.92
MARSHAL (w/o Agent-Specific Advantage Normalization)
   Tic-Tac-Toe 91.20 95.83 83.46 40.00 80.00 87.14 38.58
   Kuhn Poker 91.20 95.53 84.65 36.67 75.00 86.63 43.94
   Mini Hanabi 92.60 96.06 82.68 33.33 80.00 87.11 43.43